# Isospin asymmetry and neutron stars in V-QCD

### v1

@@ 1 Introduction
Quantum Chromodynamics is the established fundamental theory of the strong nuclear interactions. Due to asymptotic freedom, a property of non-Abelian quantum gauge theories with sufficiently little matter content, QCD is weakly coupled at high energies, where perturbative QCD (pQCD) becomes reliable. Unfortunately, the low-energy world of nuclei – protons and neutrons – is then described by QCD at strong coupling. Lattice formulation of QCD (LQCD) is a powerful numerical tool that can calculate observables on a discretized Euclidean space (a four-dimensional hypercube). Unfortunately, due to a technical problem called the sign problem deForcrand:2009zkb , LQCD struggles to converge when chemical potentials are introduced, which is needed for simulating QCD at finite densities. Neutron stars are beautifully the most compact objects in the Universe that have not collapsed into black holes and they provide us with data on mass, radius, and spin etc. – but they are at the nexus of the above-described problems: Low energies and large densities.
@@
The AdS/CFT correspondence is a powerful tool as a mathematical duality: It maps couplings, numbers and observables on one side of the duality to the other side. The fascinating property of this duality in particular, is that it also maps strong coupling to weak coupling: A strong coupled field theory on the boundary of a $d$-dimensional AdS spacetime is mapped to a weakly coupled gravitational theory in the bulk of said $d+1$ dimensional AdS. The main and strongest evidence for this duality was found by Maldacena in his seminal 1997 paper Maldacena:1997re , mapping the entire spectrum of primary operators of the superconformal field theory, i.e. $\mathcal{N}=4$ super-Yang-Mills (sYM), to AdS in five dimensions with independence from coordinates on a 5-sphere. Superconformal field theories or simply conformal field theories (CFTs) have limited usage for particle physics, with the exception of some ideas denoted unparticle physics Georgi:2007ek , since they describe field theories without reference to any scale or any size. Hence, they are not suitable for particle physics, such as QCD.
@@
Unlike the case of $\mathcal{N}=4$ sYM theory, holographic models for strong interactions, dubbed holographic QCD (HQCD), have not been derived with strong evidence for the duality to hold true and it would often be better to say they are holographic models, than dual theories to true QCD. Nevertheless, they are born with a strong-weak duality built-in and are usually constructed to assimilate the properties of strong interactions or at least main aspects thereof.
@@
Generally speaking, there are two kinds of holographic QCD models: Top-down models that are based on string theory constructions, such as the D3/D7 model Karch:2002sh , Witten-Sakai-Sugimoto model Witten:1998zw ; Sakai:2004cn , etc. They have the advantage of having a full-fledged string theory in the UV, which builds in some consistency in the models, although the UV theory is usually a web of D-branes and most likely not exactly the CFT that QCD flows to at large (infinite) energies. Another branch of HQCD models are called bottom-up as they are engineered to describe QCD purely from phenomenological considerations or even (L)QCD data. Popular models studied in recent years fall in the classes of hard-wall Polchinski:2001tt or soft-wall models Karch:2006pv .
@@
and imposing the same scaling for the gauge scaling as in the ’t Hooft limit. The “V” in the name of the model refers to the fact that the string theory inspiration of the model makes use of this limit. The V-QCD model is, however, a bottom-up model with gluon and quark degrees of freedom, where the quarks are fully backreacted to the glue background, which corresponds formally to the Veneziano limit given above. The V-QCD models parameters are determined by comparing observables to (L)QCD data at finite (and small) values of $N_{c}$ and $N_{f}$ in the end. See ref. Jarvinen:2021jbd for a recent review of the model.
Neutron stars have been considered previously in HQCD models, in particular in WSS with isospin asymmetry Kovensky:2021kzl ; Kovensky:2021wzu ; Bartolini:2023wis , in D3/D7 without isospin asymmetry Hoyos:2016zke ; BitaghsirFadafan:2019ofb ; BitaghsirFadafan:2020otb , in hard-wall without isospin asymmetry Bartolini:2022rkl , in V-QCD without isospin asymmetry Jokela:2018ers ; Chesler:2019osn ; Ecker:2019xrw ; Jokela:2020piw ; Jarvinen:2023jbr , and in V-QCD using a hybrid setup where the isospin dependence was borrowed from a field theory model Demircik:2021zll ; Tootle:2022pvd ; Ecker:2024kzs  [^1] A study of neutron star matter in an Einstein-Maxwell-dilaton-scalar theory includes isospin asymmetry, but fixes the proton fraction by hand Liu:2024efy : Isospin symmetric matter is the case where protons and neutrons are considered on equal footing as a rough approximation. Neutron stars are, however, made up of mostly neutrons as the large density allows for the baryons to be unstable (i.e. to be neutrons) because they are neutral and hence do not suffer from Coulomb repulsion. In order to take this properly into account in HQCD, isospin asymmetry must be incorporated. Although isospin asymmetric baryonic matter has been considered previously in HQCD Kovensky:2021ddl ; Bartolini:2022gdf ; Kovensky:2023mye ; Bartolini:2023wis as well as for neutron stars in the WSS model Kovensky:2021kzl ; Kovensky:2021wzu ; Bartolini:2023wis , it has not been ported to V-QCD, where the fits to LQCD infuses real-world QCD phenomenology into the model, until now.
@@
In this paper, we port for the first time isospin asymmetric baryonic matter to the precision fit-to-LQCD V-QCD holographic model of QCD, calculate the symmetry energy as well as fully nonlinearly backreacted asymmetric baryonic matter with finite isospin for the computation of nonspinning neutron stars. Using the TOV equations we compute the corresponding mass-radius curves for neutron stars and further compute the $\beta$-equilibrated neutral neutron/proton/electron/muon composition of the stars.
V-QCD has already been established as precision holographic model for quite some time. A recent and necessary development in the model was the advancement of determining the Chern-Simons (CS) term Jarvinen:2022mys . Although not important for the pure glue sector, the CS term is of utter importance for the baryon Jarvinen:2022gcc , since the stabilization of the size and position in the holographic direction depends strongly on the CS. This work on the CS is thus a necessary step as it determined the functional form and coefficients of some of the functions entering the CS term. We will, however, take a few of the parameters as free parameters in this work, see below.
@@
The paper is organized as follows. In sec. 2 we introduce the model and set our notation. In sec. 3 the homogeneous Ansatz and symmetric matter is reviewed. In sec. 4, isospin asymmetric matter is introduced into V-QCD following a generalized Ansatz previously used in the WSS and the symmetry energy is presented. In sec. 5, the possibility of a pure quark phase is considered. In sec. 6, we finally present our results on the neutron stars. We conclude with a discussion in sec. 7. Some details are delegated to appendices A and B.
@@ 2 The model
Here $M_{p}$ is the five-dimensional Planck mass, $V_{g}$ is the dilaton potential, and $h$ is the pull-back of the five-dimensional metric $g$ on the four-dimensional boundary, while $R,K$ are respectively the scalar and extrinsic curvature. The five-dimensional metric $g$ is calculated using the Ansatz
@@
Expanding the action to first nontrivial order in the gauge fields also removes an ambiguity of the DBI expression (6): As the non-Abelian gauge fields do not commute, one would need to prescribe in which order to take the trace over them. In eq. (11) this ambiguity is resolved as there are only two flavored fields in the trace.
@@
To build hybrid equations of state we need to choose a procedure to match the holographic and low-energy equations. We will present two procedures different in spirit, which we will refer to as procedures (a) and (b) from now on:
@@
For both procedures, we will use the holographic equation of state obtained with two different background choices, commonly labeled as 5b and 7a (see App. A of ref. Jokela:2020piw and App. B of ref. Ishii:2019gta for details on the definitions). These models were determined primarily by comparing to available the lattice data for the thermodynamics of QCD. There are also other fits of the V-QCD model available Amorim:2021gat ; Jarvinen:2022gcc , which also take into account data for particle spectra, but produce a less precise description of the finite temperature lattice data. For low-energy equations of state, we use the three variants from Hebeler-Lattimer-Pethick-Schwenk Hebeler:2013nza , and SLy4 Douchin:2001sv ; Haensel:1993zw . We will perform procedure (a) for three values of $n_{t}$, namely $\{1.2,1.5,1.8\}n_{S}$, for a total of 30 different equations of state, where $n_{S}$ is the phenomenological saturation density (the soft and intermediate EOS from ref. Hebeler:2013nza result in a single hybrid EOS for procedure (b)).
@@ 3 Symmetric nuclear matter
We see that $N_{B}$ vanishes for every smooth profile $H(r)$ since $H(r=0)=0$ to have a finite energy configuration Rozali:2007rx . A way to retain both a finite nonvanishing baryon number and the homogeneous field configuration is to allow the function $H(r)$ to have discontinuities Li:2015uea : The simplest choice is for it to have a discontinuity in the bulk at some finite $r=r_{c}$ (that will be determined by extremizing the action), and have $H(r>r_{c})=0$. This roughly corresponds to placing the instantons at $r=r_{c}$ before the smearing procedure. Analogously to the baryons/instantons forming a lattice in the holographic direction following “popcorn transitions” Kaplunovsky:2012gb ; Kaplunovsky:2013iza , the function $H(r)$ can in principle have multiple discontinuities (see ref. CruzRojas:2023ugm ), a refinement we do not include in the present work.
Other components of the gauge field vanish due to the absence of sources, and we are then left with two ODEs for $H(r),\widehat{a}_{0}(r)$ to be solved numerically, which we choose to do using a shooting method. A convenient choice is to shoot from the IR, since we can then fix the baryon number density $n_{B}$ and use it to impose the Dirichlet boundary condition
@@
The saturation density $n_{0}$ can be derived from the model by finding the onset of baryonic matter at which the grand potential for nuclear matter vanishes, $\Omega(n_{0})=0$. Through the rest of this work, we will keep separated the concepts of the model-derived saturation density $n_{0}$ and the phenomenological saturation density $n_{S}$. The value of the former is determined by the choice of $b_{1}$ and is subject to an intrinsic error introduced by the assumption of homogeneity, while the value of the latter is the observed one of $n_{S}=0.16\rm fm^{-3}$. When following procedure (b), the value of $b_{1}$ is chosen by requiring $n_{0}=n_{S}$, while for procedure (a) the two values in general do not coincide. Since procedure (a) is developed with the scope of matching the holographic EOS with other phenomenological ones at densities higher than saturation, then it is not a severe problem that in this case $n_{0}\neq n_{S}$.
@@ 4 Isospin asymmetry
The other components allowed in principle by the symmetries (the Abelian component for $A_{1},A_{2}$ and the missing non-Abelian components $A_{t}^{1},A_{t}^{2}$, on top of terms involving the antisymmetric symbol $\epsilon^{ij}$) are unsourced and vanishing due to the equations of motion.
@@
$4\pi^{2}M_{p}^{3}\left(\frac{e^{2A(r)}V_{f}(\phi,\tau)w(\phi)^{2}}{\sqrt{e^{2A(r)}+\kappa(\phi)\tau^{\prime 2}}}H_{3}^{\prime}\right)_{r=r_{c}}$ $=\left(\widehat{a}_{0}F_{1}(\tau)H^{2}\right)_{r=r_{c}}.$ (34)
@@ 4.1 Small isospin approximation: Symmetry energy
where we have introduced the functions $\mathcal{E}_{0}(n_{B})$, $\Lambda(n_{B})$ defined by the integrals
@@
In this analysis, we have not considered the possibility of a contribution to $S_{N}$ arising from the change of the location at which the discontinuity is placed, induced by $\mu_{I}$; that is, we determine the position in the bulk by solving $L(r_{c})=0$ at the zeroth order in $\mu_{I}$. It can be proven that the perturbation of the Lagrangian induces a perturbation of $r_{c}$ quadratic in $\mu_{I}$, so one may wonder if this perturbation allows the unperturbed Lagrangian to contribute to $S_{N}$: It can, however, be shown by using the definition of $r_{c}$ that this contribution vanishes due to the equations of motion. Thus, the isospin-induced correction to $r_{c}$ can only contribute to a higher order in $\mu_{I}$, see App. A.
@@
We have explored the symmetry energy’s ($S_{N}(n_{0})$) dependence on $b_{1}$ and $b_{3}$ at $c_{b}=1$, summarized in fig. 1. $S_{N}(n_{0})$ is an increasing function of $b_{3}$ at fixed $b_{1}$. Moreover, in the range of $b_{1}$ relevant to our matching procedures, $S_{N}(n_{0})$ is already overestimated for the flat space expression of the TCS term Casero:2007ae corresponding to $b_{3}=1$, that is, the physics of isospin asymmetry is better described for small values of $b_{3}$. This is in stark contrast with the preferred values of $b_{1}$, which governs instead the onset of nuclear matter. A minimal choice in modifying the flat space expression ($b_{1}=b_{3}=1$) for $S_{\rm TCS}$ would be to choose a single rescaling for both functions $F_{1}$, $F_{3}$, so that $b_{1}=b_{3}=b$. However, this would lead to a very large symmetry energy for both of our matching procedures (and consequently very high proton fractions for the nuclear matter in $\beta$-equilibrium). We have thus for phenomenological reasons fixed $b_{3}=1$, so that the rescaling only affects the function $F_{1}(\tau)$, to keep the number of free parameters in the model to a minimum. Without the introduction of the rescaling parameter $c_{b}$ (as in procedure (b)), this still leads to large symmetry energies, but remarkably the results for procedure (a) turns out to be close to (and often in fact compatible with) phenomenology555 It is possible that an appropriate modification of the functional form of $F_{3}$ (and $F_{1}$) may make the model reproduce phenomenological values of the symmetry energy as well as baryon onset, without the introduction of the rescaling parameter $c_{b}$. This will be beyond the scope of the present paper. .
@@
In fig. 2 is shown the symmetry energy, $S_{N}(n_{B})$, for each combination of matching procedure, potential, and parameters $c_{b}$ and $b_{1}$, which is computed as the coefficient of the expansion around saturation density (see App. B for the values of the parameters for the matching procedure (a)). The symmetry energy at saturation density resulting from matching procedure (a) lies in the range $S_{0}\in[28.7,39.4]$ MeV. Breaking down this result in the various EOSs, we find that the soft one has $S_{0}\in[28.7,37.3]$ MeV, the intermediate $S_{0}\in[31.9,37.6]$ MeV, the one matched with SLy4 $S_{0}\in[31.3,38.0]$ MeV and the stiff one $S_{0}\in[36.4,39.4]$ MeV. The lower (upper) bounds always correspond to EOS whose holographic part is built with the 5b (7a) background. Moreover, the parameter $L$ of the expansion around saturation lies in the range $L\in[51.6,71.0]$ MeV, with the upper (lower) bound corresponding again to the EOS from the 5b (7a) background. Both values are in good agreement (on the larger side) with phenomenology, a nontrivial result given the matching procedure: The rescaling with $c_{b}$ not only provides continuous EOSs, but also rescales the symmetry energy and its derivative with respect to the density to their phenomenological values. For the quantity $K_{\rm sym}$, the model predicts values $K_{\rm sym}\in[-106.1,-42.9]$ MeV.

### v2

@@ 1 Introduction
Quantum Chromodynamics is the established fundamental theory of the strong nuclear interactions. Due to asymptotic freedom, a property of non-Abelian quantum gauge theories with sufficiently little matter content, QCD is weakly coupled at high energies, where perturbative QCD (pQCD) becomes reliable. Unfortunately, the low-energy world of nuclei – protons and neutrons – is then described by QCD at strong coupling. Lattice formulation of QCD (LQCD) is a powerful numerical tool that can calculate observables on a discretized Euclidean space (a four-dimensional hypercube). Unfortunately, due to a technical problem called the sign problem deForcrand:2009zkb , LQCD struggles to converge when chemical potentials are introduced, which is needed for simulating QCD at finite densities. Neutron stars are the most compact objects in the Universe that have not collapsed into black holes and they provide us with data on mass, radius, and spin etc. – but they are at the nexus of the above-described problems: Low energies and large densities.
@@
The AdS/CFT correspondence is a powerful tool as a mathematical duality: It maps couplings, numbers and observables on one side of the duality to the other side. The fascinating property of this duality in particular, is that it also maps strong coupling to weak coupling: A strong coupled field theory on the $d$-dimensional boundary of an anti-de Sitter (AdS) spacetime is mapped to a weakly coupled gravitational theory in the bulk of said $d+1$ dimensional AdS. The main and strongest evidence for this duality was found by Maldacena in his seminal 1997 paper Maldacena:1997re , mapping the entire spectrum of primary operators of the superconformal field theory, i.e. $\mathcal{N}=4$ super-Yang-Mills (sYM), to AdS in five dimensions with independence from coordinates on a 5-sphere. Superconformal field theories or simply conformal field theories (CFTs) have limited usage for particle physics, with the exception of some ideas denoted unparticle physics Georgi:2007ek , since they describe field theories without reference to any scale or any size. Hence, they are not suitable for particle physics, such as QCD.
@@
Unlike the case of $\mathcal{N}=4$ sYM theory, holographic models for strong interactions, dubbed holographic QCD (HQCD), have not been derived with strong evidence for the duality to hold true and it would often be better to say that they are holographic models, than dual theories to true QCD. Nevertheless, they are born with a strong-weak duality built-in and are usually constructed to assimilate the properties of strong interactions or at least main aspects thereof.
@@
Generally speaking, there are two kinds of holographic QCD models: Top-down models that are based on string theory constructions, such as the D3/D7 model Karch:2002sh , the Witten-Sakai-Sugimoto (WSS) model Witten:1998zw ; Sakai:2004cn , etc. They have the advantage of having a full-fledged string theory in the UV, which builds in some consistency in the models, although the UV theory is usually a web of D-branes and most likely not exactly the CFT that QCD flows to at large (infinite) energies. Another branch of HQCD models are called bottom-up as they are engineered to describe QCD purely from phenomenological considerations or even (L)QCD data. Popular models studied in recent years fall in the classes of hard-wall Polchinski:2001tt or soft-wall models Karch:2006pv .
@@
and imposing the same scaling for the gauge coupling as in the ’t Hooft limit. The “V” in the name of the model refers to the fact that the string theory inspiration of the model makes use of this limit. The V-QCD model is, however, a bottom-up model with gluon and quark degrees of freedom, where the quarks are fully backreacted to the glue background, which corresponds formally to the Veneziano limit given above. The V-QCD models parameters are determined by comparing observables to (L)QCD data at finite (and small) values of $N_{c}$ and $N_{f}$ in the end. See ref. Jarvinen:2021jbd for a recent review of the model.
Neutron stars have been considered previously in HQCD models, in particular in WSS with isospin asymmetry Kovensky:2021kzl ; Kovensky:2021wzu ; Bartolini:2023wis , in D3/D7 without isospin asymmetry Hoyos:2016zke ; BitaghsirFadafan:2019ofb ; BitaghsirFadafan:2020otb , in hard-wall without isospin asymmetry Bartolini:2022rkl , in V-QCD without isospin asymmetry Jokela:2018ers ; Chesler:2019osn ; Ecker:2019xrw ; Jokela:2020piw ; Jarvinen:2023jbr , and in V-QCD using a hybrid setup where the isospin dependence was borrowed from a field theory model Demircik:2021zll ; Tootle:2022pvd ; Ecker:2024kzs  [^1] A study of neutron star matter in an Einstein-Maxwell-dilaton-scalar theory includes isospin asymmetry, but fixes the proton fraction by hand Liu:2024efy : Isospin symmetric matter is the case where protons and neutrons are considered on equal footing as a rough approximation. Neutron stars are, however, made up of mostly neutrons as the large density allows for the baryons to be unstable (i.e. to be neutrons) because they are neutral and hence do not suffer from Coulomb repulsion. In order to take this properly into account in HQCD, isospin asymmetry must be incorporated. Although isospin asymmetric baryonic matter has been considered previously in HQCD Kovensky:2021ddl ; Bartolini:2022gdf ; Kovensky:2023mye ; Bartolini:2023wis as well as for neutron stars in the WSS model Kovensky:2021kzl ; Kovensky:2021wzu ; Bartolini:2023wis , it has not been ported to V-QCD, where the fits to LQCD infuse real-world QCD phenomenology into the model, until now.
@@
In this paper, we port for the first time isospin asymmetric baryonic matter to the precision fit-to-LQCD V-QCD holographic model of QCD, calculate the symmetry energy as well as fully nonlinearly backreacted asymmetric baryonic matter with finite isospin for the computation of nonspinning neutron stars. Using the TOV equations we compute the corresponding mass-radius (MR) curves for neutron stars and further compute the $\beta$-equilibrated neutral neutron/proton/electron/muon composition of the stars.
V-QCD has already been established as a precision holographic model for quite some time. A recent and necessary development in the model was the advancement of determining the Chern-Simons (CS) term Jarvinen:2022mys . Although not important for the pure glue sector, the CS term is of utter importance for the baryon Jarvinen:2022gcc , since the stabilization of the size and position in the holographic direction depends strongly on the CS. This work on the CS is thus a necessary step as it determined the functional form and coefficients of some of the functions entering the CS term. We will, however, take a few of the parameters as free parameters in this work, see below.
@@
The paper is organized as follows. In sec. 2 we introduce the model and set our notation. In sec. 3 the homogeneous Ansatz and symmetric matter is reviewed. In sec. 4, isospin asymmetric matter is introduced into V-QCD following a generalized Ansatz previously used in the WSS and the symmetry energy is presented. In sec. 5, the possibility of a pure quark phase is considered. In sec. 6, we finally present our results on the neutron stars. We conclude with a discussion in sec. 7. Some details are delegated to the appendix A.
@@ 2 The model
Here $M_{p}$ is the five-dimensional Planck mass, $V_{g}$ is the dilaton potential, and $h$ is the pull-back of the five-dimensional metric $g$ on the four-dimensional boundary, while $R$ and $K$ are, respectively, the scalar and extrinsic curvature. The five-dimensional metric $g$ is calculated using the Ansatz
@@
Expanding the action to the first nontrivial order in the gauge fields also removes an ambiguity of the DBI expression (6): As the non-Abelian gauge fields do not commute, one would need to prescribe in which order to take the trace over them. In eq. (11) this ambiguity is resolved as there are only two flavored fields in the trace.
@@
To build hybrid equations of state (EOS) we need to choose a procedure to match the holographic and low-energy equations. We will present two procedures different in spirit, which we will refer to as procedures (a) and (b) from now on:
@@
For both procedures, we will use the holographic equation of state obtained with two different background choices, commonly labeled 5b and 7a (see App. A of ref. Jokela:2020piw and App. B of ref. Ishii:2019gta for details on the definitions). These models were determined primarily by comparing to available the lattice data for the thermodynamics of QCD. There are also other fits of the V-QCD model available Amorim:2021gat ; Jarvinen:2022gcc , which also take into account data for particle spectra, but produce a less precise description of the finite temperature lattice data. For low-energy equations of state, we use the three variants from Hebeler-Lattimer-Pethick-Schwenk Hebeler:2013nza , and SLy4 Douchin:2001sv ; Haensel:1993zw . We will perform procedure (a) for three values of $n_{t}$, namely $\{1.2,1.5,1.8\}n_{S}$, for a total of 30 different equations of state, where $n_{S}$ is the phenomenological saturation density (the soft and intermediate EOS from ref. Hebeler:2013nza result in a single hybrid EOS for procedure (b)).
@@ 3 Symmetric nuclear matter
We see that $N_{B}$ vanishes for every smooth profile $H(r)$ since $H(r=0)=0$ to have a finite energy configuration Rozali:2007rx . A way to retain both a finite nonvanishing baryon number and a homogeneous field configuration is to allow the function $H(r)$ to have discontinuities Li:2015uea : The simplest choice is for it to have a discontinuity in the bulk at some finite $r=r_{c}$ (that will be determined by extremizing the action), and have $H(r>r_{c})=0$. This roughly corresponds to placing the instantons at $r=r_{c}$ before the smearing procedure. Analogously to the baryons/instantons forming a lattice in the holographic direction following “popcorn transitions” Kaplunovsky:2012gb ; Kaplunovsky:2013iza , the function $H(r)$ can in principle have multiple discontinuities (see ref. CruzRojas:2023ugm ), a refinement we do not include in the present work.
We do, however, include the dynamical determination of the location of $r_{c}$, which moves toward the UV as the baryonic density increases. A possible physical interpretation of this quantity has been given in ref. Rozali:2007rx as follows: The holographic coordinate $r$ is dual to the energy scale of the dual QFT, so that the homogeneous nuclear-matter distribution in this direction would be dual to the spectrum of energies of the condensed baryons. Given that the distribution has a sharp edge at $r_{c}$, this can be interpreted as being dual to the location of a Fermi surface. In ref. Gorsky:2015pra it was observed in the context of a hard-wall model that the increase in the value of the chiral condensate leads to the movement of the baryon’s location from the IR towards the UV. Given that increasing the chiral condensate increases the energy (i.e. mass, in the case of a single instanton) of the baryon, this result is also consistent with the interpretation that $r_{c}$ should be dual to the typical energy scale of the baryon. Note that, in the model we are considering, the function $F_{1}(\tau)$ has a node at some $r_{n}$ that solves $\tau(r_{n})=\frac{1}{\sqrt{2b_{1}}}$. This implies that we can never have $r_{c}<r_{n}$, and so $r_{c}$ cannot shrink arbitrarily far from $r_{c}\sim\frac{1}{\Lambda_{\mathrm{QCD}}}$.
Other components of the gauge field vanish due to the absence of sources, and we are then left with two ordinary differential equations (ODEs) for $H(r),\widehat{a}_{0}(r)$ to be solved numerically, which we choose to do using a shooting method. A convenient choice is to shoot from the IR, since we can then fix the baryon number density $n_{B}$ and use it to impose the Dirichlet boundary condition
@@
The saturation density $n_{0}$ can be derived from the model by finding the onset of baryonic matter at which the grand potential for nuclear matter vanishes, $\Omega(n_{0})=0$. Throughout the rest of this work, we will keep separated the concepts of the model-derived saturation density $n_{0}$ and the phenomenological saturation density $n_{S}$. The value of the former is determined by the choice of $b_{1}$ and is subject to an intrinsic error introduced by the assumption of homogeneity, while the value of the latter is the observed one of $n_{S}=0.16\rm fm^{-3}$. When following procedure (b), the value of $b_{1}$ is chosen by requiring $n_{0}=n_{S}$, while for procedure (a) the two values in general do not coincide. Since procedure (a) is developed with the scope of matching the holographic EOS with other phenomenological ones at densities higher than saturation, then it is not a severe problem that in this case $n_{0}\neq n_{S}$.
@@ 4 Isospin asymmetry
The other components allowed in principle by the symmetries (the Abelian component for $A_{1},A_{2}$ and the missing non-Abelian components $A_{t}^{1},A_{t}^{2}$, on top of terms involving the antisymmetric symbol $\epsilon^{ij}$) are unsourced and vanishing due to the equations of motion. The equations of motion for the fields $A_{r}^{L},A_{r}^{R}$ are satisfied by their trivial solution, $A_{r}^{L}=A_{r}^{R}=0$, while the tachyon $\tau(r)$ is kept fixed to its background configuration (i.e. it satisfies its equation of motion in the absence of baryons), as the flavor gauge fields are taken to be in the probe approximation.
The structure of this Ansatz can be thought of as originating from the inhomogeneous configuration as follows: The holographic baryon is an instanton-like soliton, whose topological number arises from its mapping $S^{3}\rightarrow S^{3}$. The field configuration for the baryon/instanton in V-QCD is constructed in ref. Jarvinen:2022gcc , and it is the general Ansatz that can be written by imposing cylindrical symmetry together with parity and time-reversal invariance. The process of smearing the baryon on $\mathbb{R}^{3}$ then eliminates every term that is odd in $x^{i}$ (the coordinate on $\mathbb{R}^{3}$). The topological number that was originally introduced by the wrapping of $S^{3}$ on $S^{3}$ can now only be retained by the presence of a discontinuity in the fields. This picture holds both for a static instanton, in which case we obtain the Ansatz in eqs. (18)-(19), and for a rotating soliton, in which case we recover the Ansatz in eqs. (25)-(27), as appropriate. We obtain the same exact Ansatz if we start from the most general one compatible with the symmetries of the homogeneous configuration, and then carefully eliminate all unsourced fields.
@@
$4\pi^{2}M_{p}^{3}\left(\frac{e^{2A(r)}V_{f}(\phi,\tau)w(\phi)^{2}}{\sqrt{e^{2A(r)}+\kappa(\phi)\tau^{\prime 2}}}H_{3}^{\prime}\right)_{r=r_{c}}$ $=\left(\widehat{a}_{0}F_{1}(\tau)H^{2}\right)_{r=r_{c}},$ (34)
with the two equations in (31) coming from the variation of the action with respect to $\widehat{a}_{0}(r)$ and $L_{3}(r)$, respectively, and equations (32), (33) and (34) are due to the variation of the action with respect to $a_{0}(r)$, $H(r)$ and $H_{3}(r)$, respectively.
@@ 4.1 Small isospin approximation: Symmetry energy
where we have introduced the functions $\mathcal{E}_{0}(n_{B})$ and $\Lambda(n_{B})$ defined by the integrals
@@
In this analysis, we have not considered the possibility of a contribution to $S_{N}$ arising from the change of the location at which the discontinuity is placed, induced by $\mu_{I}$; that is, we determine the position in the bulk by solving $L(r_{c})=0$ at the zeroth order in $\mu_{I}$. It can be proved that the perturbation of the Lagrangian induces a perturbation of $r_{c}$ quadratic in $\mu_{I}$, so one may wonder if this perturbation allows the unperturbed Lagrangian to contribute to $S_{N}$: It can, however, be shown by using the definition of $r_{c}$ that this contribution vanishes due to the equations of motion. Thus, the isospin-induced correction to $r_{c}$ can only contribute to a higher order in $\mu_{I}$, see App. A.
@@
We have explored the symmetry energy’s ($S_{N}(n_{0})$) dependence on $b_{1}$ and $b_{3}$ at $c_{b}=1$, summarized in fig. 1. $S_{N}(n_{0})$ is an increasing function of $b_{3}$ at fixed $b_{1}$. Moreover, in the range of $b_{1}$ relevant to our matching procedures, $S_{N}(n_{0})$ is already overestimated for the flat space expression of the TCS term Casero:2007ae corresponding to $b_{3}=1$, that is, the physics of isospin asymmetry is better described for small values of $b_{3}$. This is in stark contrast with the preferred values of $b_{1}$, which governs instead the onset of nuclear matter. A minimal choice in modifying the flat space expression ($b_{1}=b_{3}=1$) for $S_{\rm TCS}$ would be to choose a single rescaling for both functions $F_{1}$, $F_{3}$, so that $b_{1}=b_{3}=b$. However, this would lead to a very large symmetry energy for both of our matching procedures (and consequently very high proton fractions for the nuclear matter in $\beta$-equilibrium). We have thus for phenomenological reasons fixed $b_{3}=1$, so that the rescaling only affects the function $F_{1}(\tau)$, to keep the number of free parameters in the model to a minimum. Without the introduction of the rescaling parameter $c_{b}$ (as in procedure (b)), this still leads to large symmetry energies, but remarkably the results for procedure (a) turns out to be close to (and in fact often compatible with) phenomenology555 It is possible that an appropriate modification of the functional form of $F_{3}$ (and $F_{1}$) may make the model reproduce phenomenological values of the symmetry energy as well as baryon onset, without the introduction of the rescaling parameter $c_{b}$. This will be beyond the scope of the present paper. .
@@
In fig. 2 is shown the symmetry energy, $S_{N}(n_{B})$, for each combination of matching procedure, potential, and parameters $c_{b}$ and $b_{1}$, which is computed as the coefficient of the expansion around saturation density (see Tab. 1 for the values of the parameters for the matching procedure (a)). The symmetry energy at saturation density resulting from matching procedure (a) lies in the range $S_{0}\in[28.7,39.4]$ MeV. Breaking down this result in the various EOSs, we find that the soft one has $S_{0}\in[28.7,37.3]$ MeV, the intermediate one has $S_{0}\in[31.9,37.6]$ MeV, the one matched with SLy4 has $S_{0}\in[31.3,38.0]$ MeV and the stiff one has $S_{0}\in[36.4,39.4]$ MeV. The lower (upper) bounds always correspond to EOS whose holographic part is built with the 5b (7a) background. Moreover, the parameter $L$ of the expansion around saturation lies in the range $L\in[51.6,71.0]$ MeV, with the upper (lower) bound corresponding again to the EOS from the 5b (7a) background. Both values are in good agreement (on the larger side) with phenomenology, a nontrivial result given the matching procedure: The rescaling with $c_{b}$ not only provides continuous EOSs, but also rescales the symmetry energy and its derivative with respect to the density to their phenomenological values. For the quantity $K_{\rm sym}$, the model predicts values $K_{\rm sym}\in[-106.1,-42.9]$ MeV.